OpenStudy (anonymous):

Giving medal if you help :) The sum of the measures of two exterior angles of a triangle is 255. What is the measure of the third exterior angle? a.)75 b.)115 c.)105 d.)95

OpenStudy (anx):

has to equal 360 degrees so 360 - 255 = 105

OpenStudy (anx):

two triangles so 180 + 180 = 360. 360 - 255 = 105

OpenStudy (anonymous):

Thank you :)